# Multi-Robot Navigation in Social Mini-Games: Definitions, Taxonomy, and Algorithms

### v2

@@ III-C5 Scalability
IV Evaluation in SMGs
We first introduce a few methods that are evaluated for the different scenarios detailed in Section II. Then we evaluate these methods using a benchmark software introduced in this work, called Social Mini-Games Library (SMGLib).
IV-A Representative SMG Solvers
Navigating from the structured domain of discrete MAPF algorithms to the domain of continuous approaches represents a significant evolution in multi-agent systems, especially within the context of SMGs. While the discrete algorithms, such as CBS [52] or ECBS [71], function within the confines of predetermined, often grid-based, spaces, continuous algorithms embrace the fluidity and intricacies of unquantized environments. This transition is akin to moving from a chessboard, with fixed squares and regimented moves, to an open field, where movements are unrestricted and can be infinitely varied. In SMGs, where nuanced interactions and dynamic adaptations are crucial, continuous algorithms provide the necessary finesse and granularity. They offer the potential for more natural and realistic agent behaviors, mirroring the seamless interactions observed in real-world social dynamics. As we venture into this continuous domain, we’ll explore strategies that harness this granularity, offering solutions that are not just mathematically efficient, but also socially coherent and intuitive. Figure 17 shows these methods graphically with multiple agents in action. In the following sections, we discuss inner-workings, strengths/limitations of some of the techniques, some of which were previously mentioned in Sec. III.
@@ IV-A1 CADRL [1]:
IV-A2 Right-Hand-Rule (RHS) [2, 3]:
@@ IV-A5 ORCA-MAPF [7, 8]:
IV-B Comparative Analysis in SMGLib
SMGLib is a simulation environment designed to connect theoretical analysis with practical applications for MRN in SMGs. SMGLib provides access to run diverse simulations for existing path planners in SMG scenarios, analyses the results, and compare with other motion planners to know the efficacy of these planners on the given benchmark scenarios (sec II-B). The library includes a variety of real-world scenarios imitating the social navigation, including both static and dynamic scenarios. The details of the usage of SMGLib are given in Appendix A.
While SMGLib is an extensible library with work in progress, in this subsection, we present the comparative analysis for the benchmarking of some MRNs in terms of their safety, smoothness, invasiveness, deadlock handling capability, and efficiency in SMGs. The analysis is done with two-agent setup in a static environment, which includes the scenarios Doorway, Intersection, and Hallway. The comparison is carried out using five evaluation metrics (sec. II-C) (Average $\triangle V$, Makespan Ratio, Path Deviation, Success-Rate, and Flow-rate), which gives an idea about the effectiveness of each MRN. Note that each scenario is simulated considering each method, but the comparison of metrics in Table I is only for the cases where the success rate is 100%.
@@ IV-B Comparative Analysis in SMGLib
TABLE I: Comparative Analysis of ORCA-MAPF, IMPC-DR, and CADRL Performance Across Environments
IV-B1 Performance of CADRL
CADRL allows agents to choose continuous actions, leading to smoother and more natural movements. This fact is visible from the least $\Delta V$ values amongst other MRNs from Table I. This flexibility helps agents navigate complex environments and avoid deadlocks by making more agile decisions. By considering the agent’s dynamic constraints, CADRL enables safer and more realistic behaviors compared to methods that only focus on velocity control. Working in a continuous action space allows agents to adapt better to different deadlock situations by selecting appropriate actions. However, one limitation is that CADRL’s optimization process may not always find the best solution, which can sometimes cause unexpected behaviors and increase the risk of deadlocks. Additionally, while CADRL supports more dynamic actions, it may require more communication and coordination among agents to prevent conflicts, which can introduce communication overhead and possible coordination issues.
IV-B2 Performance of IMPC-DR
IMPC-DR, specifically designed for deadlock avoidance, uses infinite horizon model predictive control to detect and resolve deadlocks early in a proactive manner, helping agents reach their desired goal positions. It has better performance results while determining safety, smoothness, deadlock handling, and efficiency. Typically, it has 100% success in reaching the goal without any collision and deadlock. This method has a smooth, continuous trajectory without any interruption, as depicted by the lowest Path Deviation (Table I). Due to high computational load, this method is scalable up to only 100 agents, which limits its applications for large-scale usage. In the doorway scenario, it was observed that the method has a prioritizing nature where it prioritizes the left agent to move first, then the right agent, due to the left-hand rule-induced warning band for safety.
IV-B3 Performance of ORCA-MAPF
ORCA guarantees collision avoidance by constraining agents’ velocities within a safe range. This method fails in SMGs when dealing with static and dynamic obstacles. The method does not perform well in handling deadlocks, and hence has a low success rate in various SMGs setups, failing to reach the goal safely. Further, the method has jerky navigation in more complex environments, resulting in the highest Makespan ratio for all scenarios amongst all methods from Table I. Due to the grid based map representation, there is minimal path deviation from the nominal controller, which shows the controller behaves in a non-invasive fashion.
@@ IV-B3 Performance of ORCA-MAPF
V General Trends & Open Challenges
V-A MRN in SMGs with Visual Inputs
@@
V-B MRN in SMGs with Humans
@@
V-C MRN in SMGs Using Digital Twins
@@ Appendix A SMGLib Implementation
SMGLib Design: The design of SMGLib includes a command-line interface, which makes it easy to test different scenarios for different motion-planners, quickly. The library is also extensible, thus allowing the researchers to adapt the library to their specific needs without the usual complexity. SMGLib offers interactive elements for defining the simulation space and agent characteristics. The user can select the number of agents and their state parameters, the motion-planner, type of SMG to test on, the evaluation capabilities to save the log, evaluation metrics and generate the plots. Table II gives details of these options with the supported values.
TABLE II: SMGLib Options
@@
social-orca — ORCA with social context integration
social-impc-dr — socially-aware iMPC with deadlock resolution
CADRL — Collision Avoidance with Deep RL.
Environment Scenario type: doorway, hallway, intersection.
Number of Robots Number of agents simulated per run (typically 1–4).
Agent Configuration User-defined start and goal positions within a 2D grid, $x\in[0,X_{\text{pos}}]$, $y\in[0,Y_{\text{pos}}]$.
@@ 2. Social-IMPC-DR
Figure 21: Sample log output for Robot-1 for the 3-robot doorway simulation shows the Path deviation and Average $\Delta$ V (Sec. II-C)

### v3

@@ 3.3.5 Scalability
4 Evaluation in SMGs
We first introduce a few methods that are evaluated for the different scenarios detailed in Section 2. Then we evaluate these methods using a benchmark software introduced in this work, called Social Mini-Games Library (SMGLib).
4.1 Representative SMG Solvers
Navigating from the structured domain of discrete MAPF algorithms to the domain of continuous approaches represents a significant evolution in multi-agent systems, especially within the context of SMGs. While the discrete algorithms, such as CBS [44] or ECBS [65], function within the confines of predetermined, often grid-based, spaces, continuous algorithms embrace the fluidity and intricacies of unquantized environments. This transition is akin to moving from a chessboard, with fixed squares and regimented moves, to an open field, where movements are unrestricted and can be infinitely varied. In SMGs, where nuanced interactions and dynamic adaptations are crucial, continuous algorithms provide the necessary finesse and granularity. They offer the potential for more natural and realistic agent behaviors, mirroring the seamless interactions observed in real-world social dynamics. As we venture into this continuous domain, we’ll explore strategies that harness this granularity, offering solutions that are not just mathematically efficient, but also socially coherent and intuitive. Figure 17 shows these methods graphically with multiple agents in action. In the following sections, we discuss inner-workings, strengths/limitations of some of the techniques, some of which were previously mentioned in Sec. 3.
@@ 4.1.1 CADRL [152]:
4.1.2 Right-Hand-Rule (RHS) [64, 147]:
@@ 4.1.5 ORCA-MAPF [134, 46]:
4.2 Comparative Analysis in SMGLib
SMGLib is a simulation environment designed to connect theoretical analysis with practical applications for MRN in SMGs. SMGLib provides access to run diverse simulations for existing path planners in SMG scenarios, analyses the results, and compare with other motion planners to know the efficacy of these planners on the given benchmark scenarios (sec 2.2). The library includes a variety of real-world scenarios imitating the social navigation, including both static and dynamic scenarios. The details of the usage of SMGLib are given in Appendix A.
While SMGLib is an extensible library with work in progress, in this subsection, we present the comparative analysis for the benchmarking of some MRNs in terms of their safety, smoothness, invasiveness, deadlock handling capability, and efficiency in SMGs. The analysis is done with two-agent setup in a static environment, which includes the scenarios Doorway, Intersection, and Hallway. The comparison is carried out using five evaluation metrics (sec. 2.3) (Average $\triangle V$, Makespan Ratio, Path Deviation, Success-Rate, and Flow-rate), which gives an idea about the effectiveness of each MRN. Note that each scenario is simulated considering each method, but the comparison of metrics in Table 1 is only for the cases where the success rate is 100%.
@@ 4.2 Comparative Analysis in SMGLib
Table 1: Comparative Analysis of ORCA-MAPF, IMPC-DR, and CADRL Performance Across Environments
4.2.1 Performance of CADRL
CADRL allows agents to choose continuous actions, leading to smoother and more natural movements. This fact is visible from the least $\Delta V$ values amongst other MRNs from Table 1. This flexibility helps agents navigate complex environments and avoid deadlocks by making more agile decisions. By considering the agent’s dynamic constraints, CADRL enables safer and more realistic behaviors compared to methods that only focus on velocity control. Working in a continuous action space allows agents to adapt better to different deadlock situations by selecting appropriate actions. However, one limitation is that CADRL’s optimization process may not always find the best solution, which can sometimes cause unexpected behaviors and increase the risk of deadlocks. Additionally, while CADRL supports more dynamic actions, it may require more communication and coordination among agents to prevent conflicts, which can introduce communication overhead and possible coordination issues.
4.2.2 Performance of IMPC-DR
IMPC-DR, specifically designed for deadlock avoidance, uses infinite horizon model predictive control to detect and resolve deadlocks early in a proactive manner, helping agents reach their desired goal positions. It has better performance results while determining safety, smoothness, deadlock handling, and efficiency. Typically, it has 100% success in reaching the goal without any collision and deadlock. This method has a smooth, continuous trajectory without any interruption, as depicted by the lowest Path Deviation (Table 1). Due to high computational load, this method is scalable up to only 100 agents, which limits its applications for large-scale usage. In the doorway scenario, it was observed that the method has a prioritizing nature where it prioritizes the left agent to move first, then the right agent, due to the left-hand rule-induced warning band for safety.
4.2.3 Performance of ORCA-MAPF
ORCA guarantees collision avoidance by constraining agents’ velocities within a safe range. This method fails in SMGs when dealing with static and dynamic obstacles. The method does not perform well in handling deadlocks, and hence has a low success rate in various SMGs setups, failing to reach the goal safely. Further, the method has jerky navigation in more complex environments, resulting in the highest Makespan ratio for all scenarios amongst all methods from Table 1. Due to the grid based map representation, there is minimal path deviation from the nominal controller, which shows the controller behaves in a non-invasive fashion.
@@ 4.2.3 Performance of ORCA-MAPF
5 General Trends & Open Challenges
5.1 MRN in SMGs with Visual Inputs
@@
5.2 MRN in SMGs with Humans
@@
5.3 MRN in SMGs Using Digital Twins
@@ Appendix A SMGLib Implementation
SMGLib Design: The design of SMGLib includes a command-line interface, which makes it easy to test different scenarios for different motion-planners, quickly. The library is also extensible, thus allowing the researchers to adapt the library to their specific needs without the usual complexity. SMGLib offers interactive elements for defining the simulation space and agent characteristics. The user can select the number of agents and their state parameters, the motion-planner, type of SMG to test on, the evaluation capabilities to save the log, evaluation metrics and generate the plots. Table 2 gives details of these options with the supported values.
Table 2: SMGLib Options
@@
social-orca – ORCA with social context integration
social-impc-dr – socially-aware iMPC with deadlock resolution
CADRL – Collision Avoidance with Deep RL
Environment Scenario type: doorway, hallway, intersection
Number of Robots Number of agents simulated per run (typically 1–4)
Agent Configuration User-defined start and goal positions within a 2D grid, $x\in[0,X_{\text{pos}}]$, $y\in[0,Y_{\text{pos}}]$
@@ 2. Social-IMPC-DR
Figure 21: Sample log output for Robot-1 for the 3-robot doorway simulation shows the Path deviation and Average $\Delta$ V (Sec. 2.3)